# Zhang, Liu, and Zhao

### v3

@@ 1.3 Contributions
1. We propose a PADR-based ERM method for a general class of SP problems with covariate information, where both cost and constraint functions can be nonconvex. We quantify the approximation accuracy of the PA function class and establish the nonasymptotic consistency result of our framework for unconstrained problems and asymptotic consistency result for deterministic-constrained problems without any parametric assumptions on the optimal decision rule or the latent dependency.
2. For learning PADR from the ERM problem, we develop an enhanced SMM algorithm with asymptotic convergence to composite strong d-stationarity, which is a sharper stationary concept than d-stationarity specialized to the composite DC programs with PA inner functions. We further provide nonasymptotic convergence of the algorithm to d-stationarity, which to the best of our knowledge is the first complexity result for stochastic nonsmooth composite DC programs. The enhanced SMM algorithm is also extended with exact penalization for solving the PADR-based ERM problem with constraints.
3. We test our method through numerical experiments on the Newsvendor problem and its variations with comparisons to the PO methods and other prominent DR-based methods, including the (generalized) LDR and RKHS-DR. The results demonstrate the superior performance of our method with significantly lower costs and less computation time in most instances, particularly those with high-dimensional redundant features or highly nonlinear models.
@@
The rest of this paper is organized as follows. In Section 2, we provide the theoretical consistency result of the PADR-based ERM model. We then develop the enhanced SMM algorithm for learning PADR with surrogation construction of composite PA objectives in Section 3 and present the convergence analysis in Section 4. In Section 5, we extend our framework to deterministic-constrained SP problems. In Section 6, we report numerical results. We end with conclusions in Section 7.
2 Theoretical Guarantees for PADR-based ERM
In this section, we provide theoretical guarantees of the PADR-based ERM model by first quantifying the approximation error of the PA function class to Lipschitz continuous functions and then presenting the nonasymptotic consistency result of the PADR-based ERM model via the excess risk bound. At the end of this section, we compare our theoretical results with prominent ERM models employing the LDR and RKHS-DR.
2.1 Approximation Accuracy of Piecewise Affine Function Class
@@
PA functions are known to asymptotically approximate Lipschitz continuous functions. Balazs et al. (2015) and Siahkamari et al. (2020) quantify the approximation errors of the PA function class to convex and DC functions, respectively. We provide the approximation bound of the PA function class to Lipschitz continuous functions, which to the best of our knowledge is the first quantification result for such a general case. Specifically, with $L>0$ and $M>0$, let
@@
$\mathcal{H}^{K}_{\mu}\coloneqq\left\{\left(f_{\iota}(\,\cdot\,;\theta_{\iota}):\mathcal{X}\rightarrow\mathbb{R}\right)_{\iota=1}^{d}\ \middle|\ \begin{aligned} &f_{\iota}(x;\theta_{\iota})=\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x+b^{k}_{\iota}\right\},\\ &\theta_{\iota}\coloneqq\left(\alpha_{\iota}^{k},a_{\iota}^{k},\beta_{\iota}^{k},b_{\iota}^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)},\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\,\forall\,\iota\in[d]\end{aligned}\ \right\}.$ (6)
When $K=1$, $\mathcal{H}^{1}_{\mu}$ reduces to the LDR class. We provide the approximation bound of the PA function class to Lipschitz continuous functions below with the proof given in Appendix 8.1.
Proposition 2.1 (Approximation Bound of ${\mathcal{H}}_{\mu}^{K}$)
@@
The bound $\mathcal{O}(K^{-1/p})$ provided in Proposition 2.1 is weaker than $\mathcal{O}(K^{-2/p})$ given by Balazs et al. (2015) and Siahkamari et al. (2020) in approximating convex and DC functions. Regardless, our bound is applicable to general PA-based statistical learning problems in which the underlying true model may not be convex or DC; more importantly, it renders the succeeding nonasymptotic consistency result of the PADR-based ERM model without any parametric or convex assumptions on the optimal decision rule or latent dependency.
@@ 2.2 Nonasymptotic Consistency of PADR-based ERM
We impose the following assumptions for the consistency analysis:
@@
    Samples in $S=\{(x^{s},y^{s})\}_{s=1}^{n}$ of $(X,Y)$ are independently and identically distributed.
@@
    For some $\bar{X}>0$, the support set $\mathcal{X}\subseteq\mathbb{R}^{p}$ of $X$ is a compact set with $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}=\bar{X}$.
@@
    For some $L_{\varphi}>0$, $C_{\varphi}>0$, $z_{0}\in\mathbb{R}^{d}$, and any $y\in\mathcal{Y}$, the cost function $\varphi({\cdot};{y})$ is Lipschitz continuous on $\mathbb{R}^{d}$ with modulus $L_{\varphi}$ and satisfies $\left\lvert\varphi(z_{0};y)\right\rvert\leq C_{\varphi}$.
@@
With the optimal solution to (5) denoted by $f^{*}_{S,{\mathcal{H}}_{\mu}^{K}}$, the consistency of the PADR-based ERM model is evaluated by the excess risk $R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\underset{f\in\mathcal{F}}{\inf}\,R(f)$, which is the gap between the expected cost of the optimal empirical PADR and the minimal cost. With problem (2) restricted to ${\mathcal{H}}_{\mu}^{K}$, we decompose the excess risk into the approximation and estimation errors as follows,
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)}_{\text{estimation error}}+\underbrace{\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)-\inf_{f\in\mathcal{F}}R(f)}_{\text{approximation error}}.$ (7)
@@
Under Assumptions (A1)-(A3), with $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert\leq\frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
The proof based on the Rademacher complexity theory is presented in Appendix 8.2. Combining the approximation and estimation error bounds, we obtain the excess risk bound of PADR-based ERM in Theorem 2.3 with the proof in Appendix 8.2.
@@
Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})\,\leq\,$ $\inf_{f\in\mathcal{F}}R(f)+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}\bar{X}K^{-1/p}$
@@
With $\mu=\mathcal{O}(K^{1/p})$, the complexity of the excess risk bound with respect to the sample size $n$ and parameter $K$ is
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{\widetilde{\mathcal{O}}\left(K^{\frac{p+2}{2p}}n^{-\frac{1}{2}}\right)}_{\begin{subarray}{c}\text{estimation}\\ \text{error bound}\end{subarray}}+\underbrace{\mathcal{O}\left(K^{-\frac{1}{p}}\right)}_{\begin{subarray}{c}\text{approximation}\\ \text{error bound}\end{subarray}},$ (8)
where $\widetilde{\mathcal{O}}(\cdot)$ suppresses logarithmic factors. It can be seen that the sample size $n$ only affects the estimation error, whereas $K$, the number of max-affine pieces for PA functions in ${\mathcal{H}}_{\mu}^{K}$, plays the central role of balancing the two error terms. As $K$ grows, the PA function class becomes larger, so the approximation error decreases due to the stronger approximation ability of ${\mathcal{H}}_{\mu}^{K}$, and the estimation error increases due to the more pronounced overfitting issue. By setting $K=\mathcal{O}(n^{p/(p+4)})$, we derive the optimal excess risk bound $\widetilde{\mathcal{O}}(n^{-{1}/(4+p)})$.

### v4

@@ 1.3 Contributions
1. We propose a PADR-based ERM method for a general class of SP problems with covariate information, where the cost function could be nonconvex and nondifferentiable. We quantify the approximation accuracy of the PA function class to any Lipschitz continuous function, which to the best of our knowledge is the first quantification result for such a general case. Without any parametric assumption on the optimal decision rule or the latent dependency, we establish the nonasymptotic consistency guarantee with the excess risk bound of our proposed framework, which illustrates the role of piece number in balancing the approximation and estimation errors.
2. For efficiently solving the PADR-based ERM problem, we develop an enhanced sampling-based MM (ESMM) algorithm with the natural combination of sequential sampling and the enhanced technique over the $\varepsilon$-active set. With sequential sampling, the ESMM algorithm effectively reduces computational costs by solving subproblems of small sizes in the early iterations. We show that under the proper control of increasing sample rate, ESMM achieves the asymptotic convergence in terms of composite strong d-stationarity, which is strictly stronger than d-stationarity and thus has greater potential to be a global minimizer. Numerical results show that the ESMM algorithm could speed up to 4 times faster than the basic MM method.
We also would like to highlight several distinguishing features of the proposed PADR method for contextual SP: (a) from the modeling aspect, the PADR method is suitable for contextual SP problems with nonlinear or nonconvex cost functions. All other relevant works on contextual SP with asymptotic optimality are restricted to linear programs or convex programs. For instance, the RKHS-based DR method by Bertsimas and Koduri (2021) requires convexity of the cost function for asymptotic optimality. The PADR method opens up the application to nonconvex contextual SP problems; (b) from the theoretical aspect, PADR provides a compact structure to asymptotically approximate non-parametric decision rule with minimal hyperparameter tuning. Though PA hypothesis class is not new in classical SP, we provide the first quantification of the approximation error and the excess risk bound of the data-driven PADR model for general nonconvex SP problems. Except for the LDR method, all other methods for contextual SP do not provide the excess risk bound, including the RKHS-based DR in Bertsimas and Koduri (2021) and Notz and Pibernik (2021); (c) regarding to the computational aspect, the proposed ESMM algorithm is developed for efficiently solving the PADR-based ERM problem of the large scale. Under the suitable sampling scheme, we show that the ESMM method significantly improves computational efficiency and enhances the convergence property compared to the existing MM methods in the literature; (d) moreover, the PADR method possesses several prominent numerical advantages. Numerous experiments indicate that it requires significantly fewer data points for satisfactory performance, thus is particularly suitable for cases with insufficient data sets. Additionally, it yields superior performance in terms of the test costs with both synthetic and real data sets when the feature dimension is less than 20. The nonlinearity level of the demand model has almost no effect on the performance of the PADR method, while other methods suffer significantly from the high level of nonlinearity.
@@
The rest of this paper is organized as follows. In Section 2, we provide the theoretical consistency result of the PADR-based ERM model. We develop the ESMM algorithm for learning PADR and present the convergence analysis in Section 3. In Section 4 we report numerical results. We end with conclusions in Section 5.
2 Consistency Guarantees for PADR-based ERM
In this section, we aim to derive the consistency guarantees for the PADR-based ERM problem. Considering problem (3) equipped with a class of continuous PA functions with each max-affine component having $K$ pieces and parameters $\{\theta_{\iota}\}_{\iota\in[d]}$ satisfying $\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu$ for $\iota\in[d]$, we obtain the following ERM problem:
$\min_{\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\forall\iota\in[d]}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi\left(\Big{(}\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x^{s}+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x^{s}+b^{k}_{\iota}\right\}\Big{)}_{\iota=1}^{d},y^{s}\right).$ (5)
Excess risk is commonly used to quantify the performance of the ERM model in machine learning, which can be decomposed as the approximation error for the hypothesis class and the estimation error for the empirical risk. Intuitively, when the number of pieces in the PADR class increases, the decision rule class becomes more complex, which is expected to reduce the approximation error while increasing the estimation error from the empirical risk. In this section, assuming the existence of an optimal decision rule that is Lipschitz continuous, we provide consistency guarantees of the PADR-based ERM model (5) by quantifying the approximation and estimation errors, which sheds light upon the role of the number of pieces and sample size in balancing the two errors in the excess risk. To the best of our knowledge, the approximation error bound of the PA function class established in the succeeding section is the first result for Lipschitz continuous functions in the literature. At the end of this section, we discuss the excess risk in practical implementation and compare our theoretical results with other prominent ERM models employing the LDR and RKHS-DR.
2.1 Approximation error of piecewise affine function class
@@
The approximation errors of the PA function class to convex and DC functions are quantified in Balazs et al. (2015) and Siahkamari et al. (2020) respectively. PA functions which are of “nonparametric” nature are known to asymptotically approximate Lipschitz continuous functions (Royset 2020); nevertheless, the approximation error of PA functions with finite pieces remains unsolved. For the class of Lipschitz continuous functions that allow a more general setting for SP problems of interest, we provide the approximation error bound of the PA function class, which to the best of our knowledge is the first quantification result for such a general case.
With $L>0$ and $M>0$, let
@@
$\mathcal{H}^{K}_{\mu}\coloneqq\left\{\left(f_{\iota}(\,\cdot\,,\theta_{\iota}):\mathcal{X}\rightarrow\mathbb{R}\right)_{\iota=1}^{d}\ \middle|\ \begin{aligned} &f_{\iota}(x,\theta_{\iota})=\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x+b^{k}_{\iota}\right\},\\ &\theta_{\iota}\coloneqq\left(\alpha_{\iota}^{k},a_{\iota}^{k},\beta_{\iota}^{k},b_{\iota}^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)},\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\,\forall\,\iota\in[d]\end{aligned}\ \right\}.$ (6)
When $K=1$, $\mathcal{H}^{1}_{\mu}$ reduces to the LDR class. The following proposition provides the bound for the approximation error of the PA function class to Lipschitz continuous functions.
Proposition 2.1 (Approximation error bound of ${\mathcal{H}}_{\mu}^{K}$)
@@
The idea is that for any target function $f_{0}\in\mathcal{F}_{L_{0},M_{0}}$ we construct an interpolation PA function motivated by Siahkamari et al. (2020). The difference is that with the absence of convexity, the parameters that define the interpolation functions are carefully chosen to ensure that the Lipschitz continuity modulus of the interpolation function is independent of the approximation accuracy. The constructive proof in detail is given in Online Appendix 6.1. From Proposition 2.1, the approximation error for Lipschitz continuous functions is bounded by $\mathcal{O}(K^{-1/p})$, which is weaker than the approximation error $\mathcal{O}(K^{-2/p})$ in approximating convex and DC functions by Balazs et al. (2015) and Siahkamari et al. (2020). However, it applies to general PA-based statistical learning problems in which the underlying true model is not necessarily convex or DC. Moreover, it renders the succeeding nonasymptotic consistency result of the PADR-based ERM model without any parametric or convex assumptions on the optimal decision rule or latent dependency.
@@ 2.2 Nonasymptotic consistency of PADR-based ERM
We impose the following assumptions for the nonasymptotic consistency analysis:
@@
    Samples in $\Xi=\{(x^{s},y^{s})\}_{s=1}^{n}$ of $(X,Y)$ are independently and identically distributed.
@@
    The support set $\mathcal{X}\subseteq\mathbb{R}^{p}$ of $X$ is a compact set with $\bar{X}\coloneqq\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}$.
@@
    For some $L_{\varphi}>0$, $C_{\varphi}>0$, $z_{0}\in\mathbb{R}^{d}$, and any $y\in\mathcal{Y}$, the cost function $\varphi({\cdot},{y})$ is Lipschitz continuous on $\mathbb{R}^{d}$ with modulus $L_{\varphi}$ and satisfies $\left\lvert\varphi(z_{0},y)\right\rvert\leq C_{\varphi}$.
@@
Let $f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}}$ denote the optimal solution to the PADR-based ERM model (5). The consistency of the PADR-based ERM model is evaluated by the excess risk $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\underset{f\in\mathcal{F}}{\inf}\,R(f)$, which is the gap between the expected cost of the optimal empirical PADR and the minimal cost. With $\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)$ as an intermediate problem, the excess risk can be decomposed into the approximation and estimation errors as follows,
$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)}_{\text{estimation error}}+\underbrace{\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)-\inf_{f\in\mathcal{F}}R(f)}_{\text{approximation error}}.$ (7)
@@
Under Assumptions (A1)-(A3), with $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq\frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, with probability at least $1-\delta$ over the draw of $\Xi$, we have
$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})\,\leq\,$ $\inf_{f\in\mathcal{F}}R(f)+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}\bar{X}K^{-1/p}$
@@
The above theorem implies that with $\mu=\mathcal{O}(K^{1/p})$, the complexity of the excess risk bound with respect to the sample size $n$ and parameter $K$ is
$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{\widetilde{\mathcal{O}}\left(K^{\frac{p+2}{2p}}n^{-\frac{1}{2}}\right)}_{\begin{subarray}{c}\text{estimation}\\ \text{error bound}\end{subarray}}+\underbrace{\mathcal{O}\left(K^{-\frac{1}{p}}\right)}_{\begin{subarray}{c}\text{approximation}\\ \text{error bound}\end{subarray}},$ (8)
where $\widetilde{\mathcal{O}}(\cdot)$ suppresses logarithmic factors. It can be seen that the sample size $n$ only affects the estimation error, whereas $K$, the number of max-affine pieces for PA functions in ${\mathcal{H}}_{\mu}^{K}$, plays the central role of balancing the two error terms. As $K$ grows, the PA function class becomes larger, so the approximation error decreases due to the stronger approximation ability of ${\mathcal{H}}_{\mu}^{K}$, and the estimation error increases due to the more pronounced overfitting issue. With an appropriate choice of $K$ that balances these two errors, we can derive the optimal excess risk bound. Furthermore, with the existence of a convex or DC optimal DR, this bound can be improved based on the approximation bounds $\mathcal{O}(K^{-2/p})$ of ${\mathcal{H}}_{\mu}^{K}$ from Balazs et al. (2015) and Siahkamari et al. (2020). The extensions are summarized in the following corollary.
Corollary 2.4
Under the settings in Theorem 2.3, with $K=\mathcal{O}(n^{p/(p+4)})$, the optimal excess risk bound is $\widetilde{\mathcal{O}}(n^{-{1}/(4+p)})$. If there exists a convex or DC optimal decision rule to problem (2), the optimal excess risk bound is $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$ with $K=\mathcal{O}(n^{p/(p+6)})$.
We make a remark that Proposition 2.1 provides a bound for the worst-case approximation error and thus the excess risk bound (8) could be conservative in practice. In the practical implementation of PADR in Section 4, we observe that a much more moderate number of $K$ is able to produce satisfactory results including the two-stage product placement problem and Newsvendor problems with real data. This is because the (near) optimal decision rule may possess some underlying structural property besides Lipschitz continuity so that it could be approximated well by the PADR with a moderate number of pieces, for which a smaller estimation error and excess risk could be achieved.
The excess risk bound (8) illustrates the role of $K$ in balancing the two errors, which is further evidenced by numerical experiments summarized in the table below. Table 2.2 reports the validation and training costs provided by the PADR-based ERM method in solving a Newsvendor problem with the sine plus max-affine demand model (19) (details are given in Section 4.1), under multiple choices of $K_{1}$ and $K_{2}$ which represent the number of pieces for the two max-affine components in the PADR respectively. The validation cost, which is the empirical cost of the obtained PADR solution over the validation data set of size 1000, can be regarded as the sample average approximation of $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})$. Noticing that $\inf_{f\in\mathcal{F}}R(f)$ is irrelevant to the PA hypothesis class, the performance of excess risk could be revealed through the pattern of validation costs in the table below. For $K_{1}$ and $K_{2}$ being small, the validation cost decreases along with increasing values of $K_{1}$ and $K_{2}$ which may be credited to the decreasing approximation error. For $K_{1}$ and $K_{2}$ being larger than 4, the validation cost slightly increases, suggesting that as the PADR model becomes more complex, the increasing estimation error gradually dominates the decreasing approximation error. Hence, the best validation costs shall be achieved with moderate values of $K_{1}$ and $K_{2}$, which are usually considerably smaller than the ones implied by the theoretical results.
[…]
Validation and training costs for basic Newsvendor with demand model (19) $K_{1}\backslash K_{2}$ 1 2 3 4 5 6 1 24.06 (24.09) 23.62 (23.57) 23.60 (23.63) 23.64 (23.55) 23.72 (23.66) 23.64 (23.62) 2 8.68 (8.73) 7.01 (7.09) 7.01 (7.09) 7.02 (7.08) 7.00 (7.10) 7.02 (7.09) 3 7.82 (7.95) 5.41 (5.53) 5.45 (5.51) 5.44 (5.51) 5.43 (5.49) 5.42 (5.49) 4 7.29 (7.47) 4.26 (4.27) 4.25 (4.22) 4.28 (4.29) 4.29 (4.21) 4.22 (4.23) 5 7.26 (7.46) 4.26 (4.27) 4.25 (4.24) 4.23 (4.21) 4.25 (4.21) 4.27 (4.22) 6 7.28 (7.47) 4.27 (4.24) 4.17 (4.23) 4.16 (4.23) 4.24 (4.27) 4.22 (4.20) The PADR is obtained by Algorithm 1 for solving the PADR-based ERM model. Numbers listed in brackets represent training costs, and the lowest costs are highlighted in bold and red text.